# Convergent Privacy Framework with Contractive GNN Layers for Multi-hop Aggregations

### v1

@@ Appendix C Proofs for results in Section 5
[…]
@@
C.2. Proof of results in Section 5.2
@@
The proof is reproduced from the original proof in Bok et al. (2024), we include it here for self-containedness and to clarify that the contractive condition (Lipschitz constant bounded by $\gamma<1$) is only required for iteration steps $k\geq 2$.
@@
The key idea in the proof, following Bok et al. (2024), is to construct an auxiliary interpolating sequence $\{\widetilde{{\bm{X}}}^{(k)}\}_{k=0}^{K}$ between the two CNI processes. For each step $k$, we define:
(11) $\widetilde{{\bm{X}}}^{(k+1)}=\Pi_{{\mathcal{K}}}\bigl{(}\lambda_{k+1}\phi_{k+1}({\bm{X}}^{(k)})+(1-\lambda_{k+1})\phi^{\prime}_{k+1}(\widetilde{{\bm{X}}}^{(k)})+{\bm{Z}}^{(k+1)}\bigr{)},$
@@
We now recall a lemma from Bok et al. (2024) that establishes the trade-off function bound for one step of interpolation.
Lemma 0 ((Bok et al., 2024, Lemma 3.2)).
@@
$\|{\bm{X}}^{(k)}-\widetilde{{\bm{X}}}^{(k)}\|$ $\leq\|\Pi_{{\mathcal{K}}}(\phi_{k}({\bm{X}}^{(k-1)})+{\bm{Z}}^{(k)})$
$-\Pi_{{\mathcal{K}}}(\lambda_{k}\phi_{k}({\bm{X}}^{(k-1)})+(1-\lambda_{k})\phi^{\prime}_{k}(\widetilde{{\bm{X}}}^{(k-1)})+{\bm{Z}}^{(k)})\|$
@@
where the second inequality follows from the fact that $\Pi_{{\mathcal{K}}}$ is a projection onto a convex set and hence has Lipschitz constant $1$ ((Bok et al., 2024, Lemma 2.9)). This concludes the proof of the claim by induction. We let $a_{k+1}=\lambda_{k+1}(c_{k+1}z_{k}+s)$ for $k=0,\dots,K-1$. In particular, $a_{1}=\lambda_{1}s$ regardless of the value $c_{1}$ since it is multiplied with $z_{0}=0$. This is the main our observation that the Lipchitz constant of the first iteration does not affect the constant $a_{1}$ and consequently the privacy guarantee.
@@
Lastly, by the assumption that $\phi_{k}$ and $\phi^{\prime}_{k}$ are $\gamma$-Lipschitz for $k=2,\dots,K$, we can let $c_{k}=\gamma$ for $k=2,\dots,K$. We set $\lambda_{k}=\frac{\gamma^{K-k}(1-\gamma^{2})}{1-\gamma^{K-k+2}-\gamma^{k}+\gamma^{K}}$ for $k=1,\dots,K$ as computed in Bok et al. (2024, Lemma C.6). Then $z_{k}=\frac{(1-\gamma^{k})(1-\gamma^{K-k})}{(1+\gamma^{K})(1-\gamma)}s$, and consequently, $a_{k}=\frac{\gamma^{K-k}(1+\gamma)}{1+\gamma^{k}}s$. In this case, $\sum_{k=1}^{K}a_{k}^{2}=\frac{(1-\gamma^{K})(1+\gamma)}{(1+\gamma^{K})(1-\gamma)}s^{2}$. This concludes the proof of the theorem. ∎
@@
C.3. Proof of results in Section 5.3
@@
$\leq C_{\mathsf{L}}\|\alpha_{1}(\hat{{\bm{A}}}{\bm{Y}}^{(k-1)}-\hat{{\bm{A}}}{\bm{Y}}^{\prime(k-1)})+\alpha_{2}(\mathsf{Mean}({\bm{Y}}^{(k-1)})-\mathsf{Mean}({\bm{Y}}^{\prime(k-1)}))\|$
$\leq C_{\mathsf{L}}(\alpha_{1}+\alpha_{2})\|{\bm{Y}}^{(k-1)}-{\bm{Y}}^{\prime(k-1)}\|=C_{\mathsf{L}}\|{\bm{Y}}^{(k-1)}-{\bm{Y}}^{\prime(k-1)}\|,$
@@
The difference between the two outputs is given by the aggregation of $\hat{{\bm{A}}}$ and $\hat{{\bm{A}}}^{\prime}$. Then the edge-level sensitivity $\Delta_{e}(\mathsf{CGL})$ is amount to bound $\|\hat{{\bm{A}}}{\bm{X}}^{(k)}-\hat{{\bm{A}}}^{\prime}{\bm{X}}^{\prime(k)}\|_{F}$. Since only one edge is added, the difference between $\hat{{\bm{A}}}$ and $\hat{{\bm{A}}}^{\prime}$ is only on the row corresponding to $u$ and $v$.
@@
(12) $(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}$ $=\frac{1}{d_{u}+1}{\bm{X}}^{(k)}_{u}+\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+1}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}$
@@
(13) $(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}$ $=\frac{1}{d_{u}+2}{\bm{X}}^{(k)}_{u}+\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}$
@@
$+\quad\left\|\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+1}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}-\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}\right\|_{2}$
@@
$\leq\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{D_{\min}}{\sqrt{D_{\min}+1}}(\frac{1}{\sqrt{D_{\min}+1}}-\frac{1}{\sqrt{D_{\min}+2}})$
$+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}$
@@
(14) $C(D_{\min})=\begin{cases}\frac{D_{\min}}{\sqrt{D_{\min}+1}}-\frac{D_{\min}}{\sqrt{D_{\min}+2}}&D_{\min}>3\\ (\frac{3}{\sqrt{4}}-\frac{3}{\sqrt{5}})&1\leq D_{\min}\leq 3\\ \end{cases}$
Therefore, the affect of modifying an edge on a single node $u$ of $\hat{{\bm{A}}}{\bm{X}}^{(k)}$ is bounded by
(15) $\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$
$\leq\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}$
@@
$\leq\alpha_{1}C_{\mathsf{L}}\sqrt{\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}^{2}+\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{v}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{v}\|_{2}^{2}}$
$\leq\sqrt{2}\alpha_{1}C_{\mathsf{L}}\left(\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}\right)$
@@
$+\quad\left\|\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+1}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}-\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}\right\|_{2}$
@@
$\leq\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}$
@@
$\leq\sum_{u\in{\bm{N}}_{v}}\left(\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}\right)$
$\leq|d_{v}^{\prime}|\left(\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}\right)$
$\leq\frac{\sqrt{d_{v}^{\prime}}}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})\sqrt{d_{v}^{\prime}}}{\sqrt{D_{\min}+1}}+\frac{\sqrt{d_{v}^{\prime}}}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}$
@@
$\leq 1+\alpha_{1}C_{\mathsf{L}}\left(\frac{\sqrt{D_{\max}}}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})\sqrt{D_{\max}}}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}}\right)$
@@ Appendix D Experimental Setup
D.1. Datasets
Table 5 lists statistics of these datasets, including Amazon co-purchase networks (Photo, Computers (Shchur et al., 2018)), social network (Facebook (Traud et al., 2011)), and citation networks (Cora, PubMed (Sen et al., 2008)). These datasets are also widely adopted as benchmark datasets to evaluate various GNNs (Gu et al., 2020; Sajadmanesh et al., 2023; Chien et al., 2023; Yao et al., 2023). In addition, we use the synthetic graph generation algorithm developed in IGNN (Gu et al., 2020), to produce chain-structured datasets for convenient and generalized verification of multi-hopo aggregations. The chain-structured dataset can be configured with various number of nodes per chain, number of chains per class, and the number of class. In Table 6, the statistics of the four synthesized chain-structured datasets, namely Chain-S, Chain-M, Chain-L, and Chain-X  are shown. We unify the feature dimension to $5$ and focus on the structure of chain-structured dataset, and we configure the number of nodes to $\{8,10,15\}$ per chain, and the number of chains to $\{3,5\}$ per class.
Table 5. Standard Graph Dataset Statistics. “#Tra” and “#Test” are the ratios of the total nodes used for training and testing.
@@ D.1. Datasets
Facebook $26,406$ $2,117,924$ $501$ $6$ $10\%$ $20\%$
PubMed $19,717$ $88,648$ $500$ $3$ $10\%$ $20\%$
Cora $2,708$ $10,556$ $1,433$ $7$ $10\%$ $20\%$
Photo $7,535$ $238,162$ $745$ $8$ $10\%$ $20\%$
Table 6. Chain-structured Datasets Statistics. Each chain has the same number of nodes and “Node” is the sum of nodes across chains.
@@
Chain-S $48$ $6$ $5$ $2$ $8$ $32$
Chain-M $60$ $6$ $5$ $2$ $10$ $40$
Chain-L $90$ $6$ $5$ $2$ $15$ $60$
Chain-X $150$ $10$ $5$ $2$ $25$ $100$
D.2. Baseline Methods
Edge-level private algorithms.

### v2

@@ A-A Preliminaries: Perturbed Message Passing
By the Gaussian mechanism guarantee under Rényi differential privacy (RDP) [csf/mironov2017renyi, Corollary 3], the map $X\mapsto\mathrm{MP}_{{\bm{G}}}(X)+Z$ is $\bigl(\alpha,\alpha\,\Delta^{2}(\mathrm{MP})/(2\sigma^{2})\bigr)\text{-RDP}.$ The post-processing theorem then ensures that projecting to $\mathcal{K}$} does not increase the privacy cost. Converting from RDP to $(\epsilon,\delta)\text{-DP}$ [csf/mironov2017renyi, Proposition 3] completes the proof. ∎
@@
A-B Proof of results in Section V-A
@@
The proof is reproduced from the original proof in [icml/BokShifted24], we include it here for self-containedness and to clarify that the contractive condition (Lipschitz constant bounded by $\gamma<1$) is only required for iteration steps $k\geq 2$.
@@
The key idea in the proof, following [icml/BokShifted24], is to construct an auxiliary interpolating sequence $\{\widetilde{{\bm{X}}}^{(k)}\}_{k=0}^{K}$ between the two CNI processes. For each step $k$, we define:
$\widetilde{{\bm{X}}}^{(k+1)}=$ $\Pi_{{\mathcal{K}}}\bigl(\lambda_{k+1}\phi_{k+1}({\bm{X}}^{(k)})$ (11)
$\quad+(1-\lambda_{k+1})\phi^{\prime}_{k+1}(\widetilde{{\bm{X}}}^{(k)})+{\bm{Z}}^{(k+1)}\bigr),$
@@
We now recall a lemma from [icml/BokShifted24] that establishes the trade-off function bound for one step of interpolation.
Lemma 4 ([icml/BokShifted24, Lemma 3.2]).
@@
$\|{\bm{X}}^{(k)}-\widetilde{{\bm{X}}}^{(k)}\|X\leq\|\Pi_{{\mathcal{K}}}(\phi_{k}({\bm{X}}^{(k-1)})+{\bm{Z}}^{(k)})$
$\quad-\Pi_{{\mathcal{K}}}(\lambda_{k}\phi_{k}({\bm{X}}^{(k-1)})+(1-\lambda_{k})\phi^{\prime}_{k}(\widetilde{{\bm{X}}}^{(k-1)})+{\bm{Z}}^{(k)})\|$
@@
where the second inequality follows from the fact that $\Pi_{{\mathcal{K}}}$ is a projection onto a convex set and hence has Lipschitz constant $1$ ([icml/BokShifted24, Lemma 2.9]). This concludes the proof of the claim by induction. We let $a_{k+1}=\lambda_{k+1}(c_{k+1}z_{k}+s)$ for $k=0,\dots,K-1$. In particular, $a_{1}=\lambda_{1}s$ regardless of the value $c_{1}$ since it is multiplied with $z_{0}=0$. This is the main our observation that the Lipchitz constant of the first iteration does not affect the constant $a_{1}$ and consequently the privacy guarantee.
@@
Lastly, by the assumption that $\phi_{k}$ and $\phi^{\prime}_{k}$ are $\gamma$-Lipschitz for $k=2,\dots,K$, we can let $c_{k}=\gamma$ for $k=2,\dots,K$. We set $\lambda_{k}=\frac{\gamma^{K-k}(1-\gamma^{2})}{1-\gamma^{K-k+2}-\gamma^{k}+\gamma^{K}}$ for $k=1,\dots,K$ as computed in [icml/BokShifted24, Lemma C.6]. Then $z_{k}=\frac{(1-\gamma^{k})(1-\gamma^{K-k})}{(1+\gamma^{K})(1-\gamma)}s$, and consequently, $a_{k}=\frac{\gamma^{K-k}(1+\gamma)}{1+\gamma^{k}}s$. In this case, $\sum_{k=1}^{K}a_{k}^{2}=\frac{(1-\gamma^{K})(1+\gamma)}{(1+\gamma^{K})(1-\gamma)}s^{2}$. This concludes the proof of the theorem. ∎
@@
A-C Proof of results in Section V-B
@@
$\leq C_{\mathsf{L}}\|\alpha_{1}(\hat{{\bm{A}}}{\bm{Y}}^{(k-1)}-\hat{{\bm{A}}}{\bm{Y}}^{\prime(k-1)})$
$\quad\quad+\alpha_{2}(\mathsf{Mean}({\bm{Y}}^{(k-1)})-\mathsf{Mean}({\bm{Y}}^{\prime(k-1)}))\|$
$\leq C_{\mathsf{L}}(\alpha_{1}+\alpha_{2})\|{\bm{Y}}^{(k-1)}-{\bm{Y}}^{\prime(k-1)}\|$
$=C_{\mathsf{L}}\|{\bm{Y}}^{(k-1)}-{\bm{Y}}^{\prime(k-1)}\|,$
@@
The difference between the two outputs is given by the aggregation of $\hat{{\bm{A}}}$ and $\hat{{\bm{A}}}^{\prime}$. Then the edge-level sensitivity $\Delta_{e}(\mathsf{CGL})$ is the amount to bound $\|\hat{{\bm{A}}}{\bm{X}}^{(k)}-\hat{{\bm{A}}}^{\prime}{\bm{X}}^{\prime(k)}\|_{F}$. Since only one edge is added, the difference between $\hat{{\bm{A}}}$ and $\hat{{\bm{A}}}^{\prime}$ is only on the row corresponding to $u$ and $v$.
@@
$(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}$ $=\frac{1}{d_{u}+1}{\bm{X}}^{(k)}_{u}+\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+1}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}$ (12)
@@
$(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}$ $=\frac{1}{d_{u}+2}{\bm{X}}^{(k)}_{u}+\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}$ (13)
@@
$+\left\|\sum_{w\in{\bm{N}}_{u}}\hskip-5.69054pt\frac{1}{\sqrt{d_{u}+1}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}-\hskip-5.69054pt\sum_{w\in{\bm{N}}_{u}}\hskip-5.69054pt\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}\right\|_{2}$
@@
$\leq\frac{1}{(D_{\min}+1)(D_{\min}+2)}$
$\quad+\frac{D_{\min}}{\sqrt{D_{\min}+1}}(\frac{1}{\sqrt{D_{\min}+1}}-\frac{1}{\sqrt{D_{\min}+2}})$
$\quad+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}$
@@
[…]
Therefore, the effect of modifying an edge on a single node $u$ of $\hat{{\bm{A}}}{\bm{X}}^{(k)}$ is bounded by
$\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$ (15)
$\leq\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}$
$\quad+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}$
@@
[…]
$\leq\sqrt{2}\alpha_{1}C_{\mathsf{L}}\left(\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}\right$.
$\qquad\left.+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}\right)$
@@
$+\left\|\sum_{w\in{\bm{N}}_{u}}\hskip-5.69054pt\frac{1}{\sqrt{d_{u}+1}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}-\sum_{w\in{\bm{N}}_{u}}\hskip-5.69054pt\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}\right\|_{2}$
@@
$\leq\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}$
$+\frac{1}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}$
@@
$\leq\sum_{u\in{\bm{N}}_{v}}\left(\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}\right$.
$\qquad\left.+\frac{1}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}\right)$
$\leq|d_{v}^{\prime}|\left(\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}\right$.
$\qquad\left.+\frac{1}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}\right)$
$\leq\frac{\sqrt{d_{v}^{\prime}}}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})\sqrt{d_{v}^{\prime}}}{\sqrt{D_{\min}+1}}$
$\qquad+\frac{\sqrt{d_{v}^{\prime}}}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}$
@@
$\leq 1+\alpha_{1}C_{\mathsf{L}}\left(\frac{\sqrt{D_{\max}}}{(D_{\min}+1)(D_{\min}+2)}\right$.
$\qquad\left.+\frac{C(D_{\min})\sqrt{D_{\max}}}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}}\right)$
@@ A-D Experimental Setup
A-D1 Datasets
Table V lists statistics of these datasets, including Amazon co-purchase networks (Photo, Computers [corr/abs-1811-05868]), social network (Facebook [corr/abs-1102-2166]), and citation networks (Cora, PubMed [sen2008collective]). These datasets are also widely adopted as benchmark datasets to evaluate various GNNs [nips/GuC0SG20, uss/sajadmanesh2023gap, nips/chien2023differentially, nips/YaoJRJ23]. In addition, we use the synthetic graph generation algorithm developed in IGNN [nips/GuC0SG20], to produce chain-structured datasets for convenient and generalized verification of multi-hop aggregations. The chain-structured dataset can be configured with various number of nodes per chain, number of chains per class, and the number of classes. In Table VI, the statistics of the four synthesized chain-structured datasets, namely Chain-S, Chain-M, Chain-L, and Chain-X  are shown. We unify the feature dimension to $5$ and focus on the structure of chain-structured dataset, and we configure the number of nodes to $\{8,10,15\}$ per chain, and the number of chains to $\{3,5\}$ per class.
TABLE V: Standard Graph Dataset Statistics. “#Tra” and “#Test” are the ratios of the total nodes used for training and testing.
@@ A-D1 Datasets
Facebook $26,406$ $2,117,924$ $501$ 66 $10\%$ $20\%$
PubMed $19,717$ $88,648$ $500$ 33 $10\%$ $20\%$
Cora $2,708$ $10,556$ $1,433$ 77 $10\%$ $20\%$
Photo $7,535$ $238,162$ $745$ 88 $10\%$ $20\%$
TABLE VI: Chain-structured Datasets Statistics. Each chain has the same number of nodes and “Node” is the sum of nodes across chains.
@@
Chain-S $48$ 66 55 22 88 $32$
Chain-M $60$ 66 55 22 $10$ $40$
Chain-L $90$ 66 55 22 $15$ $60$
Chain-X $150$ $10$ 55 22 $25$ $100$
A-D2 Hardware and software
Our experiments have been conducted on the Ubuntu 20.04.2 LTS server, with AMD Ryzen Threadripper 3970X 32-core CPUs of 256 GB CPU memory and NVIDIA GeForce RTX 3090 of 24GB memory. $\mathsf{CARIBOU}$ equipped with edge- and node-level privacy has been implemented in PyTorch using PyTorch-Geometric (PyG) [corr/abs-1903-02428] framework. DP implementation adopts the autodp library [AutoDP2025], which includes analytical moments accountant and private training with DP-SGD [ccs/AbadiCGMMT016].
A-D3 Baseline algorithms